OpenStudy (anonymous):

Solve for y:2/3 y + 4 = 6 y = 3 y = 7 y = -3 y = -7

OpenStudy (anonymous):

2/3 y + 4 = 6 -4 -4 2/3y=2 3(2/3y)=3(2) 2y=6 y=3

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Your Welcome!